# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

### v2

@@ 3.2 Data Analysis and Motivations
Spatial Relevance. TkQ congEfficientRetrievalTopk2009 ; liu2023effectiveness typically computes spatial relevance by $1-SDist(q.loc,o.loc)$. The implicit assumption behind this linear function is that the user’s geographical preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and we found it does not hold on our real-life datasets. In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$. This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario. This motivates us to design a new spatial relevance module to better fit the real-world scenario. An intuitive solution would be to employ an exponential function. However, our experiments indicate that this method is even less effective compared to the linear function (as to be shown in Section 5.5).
@@
Motivations of LIST. As discussed in the last paragraph, incorporating deep textual relevance into spatial keyword queries has presented significant efficiency challenges. However, as discussed in Section 2, the first category of deep text relevance model, namely word interaction based deep relevance models, such as DrW liu2023effectiveness , which rely on word interaction modules and word embeddings to compute textual relevance, are not efficient for online computation and retrieval by a specifically designed index. Therefore, developing a new lightweight yet effective relevance model and a new index to accelerate the top-k search process for spatial keyword queries based on the proposed deep relevance model is in demand. For the first objective, we develop a lightweight yet effective deep relevance model for embedding based spatial keyword queries.
For the second objective, existing geo-textual indexes DBLP:journals/vldb/ChenCCJ21 developed for spatial keyword queries are designed based on traditional retrieval models such as BM25 and cannot be used to handle embedding based spatial keyword queries. Existing ANNS indexes, which are discussed in Section 2, do not consider spatial factors. Directly employing existing ANNS indexes for spatial keyword queries will result in severe degradation of effectiveness (shown in Section 5.2). Simply adapting existing ANNS indexes, such as the IVF index, to incorporate spatial factors will cause even more severe effectiveness degradation, as it is hard to manually set a weight to balance spatial and textual similarities (as to be shown in Section 5.2). It is still an open problem to design a new ANNS index to support embedding based spatial keyword queries.
3.3 Problem Statement
In this study, we aim to (1) design a lightweight yet effective relevance model, and (2) develop an index to expedite the top-k search process of the proposed relevance model.
@@ 4 Proposed Retriever (LIST)
Figure 2: The three phases of our retriever LIST: the training, indexing, and query phase. LIST consists of a relevance model (shown in yellow) and an index (shown in green).
@@ 4.1 Overview
LIST is featured with a new relevance model and a novel machine learning based ANNS index. Notably, our proposed index is also applicable to any other textual relevance model that transforms query keywords and object text descriptions into two separate embeddings.
The workflow of LIST is illustrated in Figure 2, which has three phases: the training, indexing, and query phase. During the training phase, we first train our relevance model, and then train our index. During the indexing phase, each object is partitioned into one cluster. During the query phase, the trained index routes each query to either a single cluster or a subset of clusters that have the highest probabilities. Within these clusters, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
4.2 The Proposed Relevance Model
Textual Relevance Learning. Inspired by the dual-encoder model’s success in document retrieval DBLP:conf/emnlp/KarpukhinOMLWEC20 , we employ a dual-encoder module to encode query keywords and object description into two seperate embeddings, and calculate textual relevance by the inner product between the two embeddings. Compared with previous relevance models liu2023effectiveness that rely on word embeddings and complex word interaction functions discussed in Section 2, the dual encoder module is both efficient and effective.
The dual-encoder module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, which are pre-trained language models like Bert devlin2018bert . The encoder takes the textual content of the query or object as input, captures interactions between words by a transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional embedding. This process is formulated as:
@@ 4.2 The Proposed Relevance Model
As discovered in Section 3.2, the user’s geographic preferences for geo-textual objects do not follow a linear pattern. Here, we consider that the user’s geographic preference for geo-textual objects increases as the distance decreases, i.e., users tend to visit nearby objects, and exhibits a stepwise decline as the distance increases. To explain this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks. If the nearest Starbucks is very close, s/he would go and buy coffee. If the nearest Starbucks is a little far, s/he may hesitate and her/his intention of visiting the Starbucks would decrease. If the nearest Starbucks is very far away, s/he would give up this idea. Therefore, the new spatial relevance module is designed to be a learnable monotonically increasing step function, to align with this real-world characteristic.
@@
Adaptive Weight Learning. The recent study liu2023effectiveness has shown the importance of weight learning in improving ranking effectiveness. To enhance ranking effectiveness, we propose an adaptive weight learning module, which aims to assign adaptive weights to textual and spatial relevance based on the query keywords. For instance, in the case of detailed query keywords, such as ‘White House, 1600 Pennsylvania Avenue NW, Washington’, giving greater weight to textual relevance better fits the real-world scenario. Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine the weights based on the embedding of query keywords $q.emb$, which is formulated as follows:
@@
Training Strategy. To train our relevance model, we employ the contrastive learning strategy. Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section 5.1). As for the negative (irrelevant) geo-textual objects, following previous study liu2023effectiveness , in each training epoch, we randomly pick $b$ hard negative samples from the dataset. We optimize the loss function as the negative log-likelihood of the positive object:
@@ 4.3 The Proposed Index
Feature Construction. We utilize geo-location and textual embedding to construct a consistent input representation for both objects and queries, which is illustrated in Figure 3. The textual embedding $emb\in\mathbb{R}^{d}$ is a $d$-dimensional embedding, which is converted from textual content $o.doc$ or $q.doc$ by the trained dual encoder module. The geo-location $loc=\langle lat,lon\rangle$ are transformed into the following features $\langle\hat{lat},\hat{lon}\rangle$ as below:
@@
Here $x$ is the input feature vector for the cluster classifier, i.e., the MLP.
@@
Insertion and Deletion Policy. When a new object comes, we convert it into an embedding using the trained relevance model and then assign it to specific clusters using the trained index. When an object is deleted, we simply remove it from the corresponding cluster. Note that most existing ANNS indexes are static luoSurveyDeepHashing2023 ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 . When inserted data significantly differs in distribution from the existing data or when insertions occur frequently, periodically rebuilding the index is necessary to maintain high accuracy. Our index follows this approach. Apparently, the time costs of insertion and deletion operations are negligible.
@@ 18       $Res_{q}\leftarrow\text{argTop-k}_{o\in C_{i}}ST(q,o)$;
Procedures of LIST. The detailed procedures of LIST are summarized at Algorithm 1. In the training phase, we train the proposed relevance model and index (lines 2-4). After that, in the indexing phase, we assign all objects to their corresponding clusters as inverted files (lines 6-10). In the query phase, given a new query $q$, we extract its features and then route it to a cluster $C_{i}$ by the index (lines 13-16). Consequently, we calculate the relevance score between $q$ and each object $o$ in that cluster by the relevance model, and the top-$k$ objects with the highest scores are returned to answer query $q$ (lines 17-18).
Complexity Analysis. Now we analyze the time and space complexity of LIST. Assuming the dimensionality of the embedding is $d$, the total number of objects is $n$, and the number of layers in the MLP of the index is $l$. The embeddings $o.emb$ of objects are generated in advance.
For a given query, the time complexity of LIST is $O((l-2)d^{2}+dc+\frac{n}{c}(d+2)+d)$. $O((l-2)d^{2}+dc)$ is the time complexity for our index (lines 15), which is the inference computation cost of Equation 11. $O(\frac{n}{c}(d+2)+d)$ is the time cost of our relevance model (line 17). $\frac{n}{c}$ denotes the number of objects to be checked by the relevance model, which is approximately $\frac{1}{c}$ of the entire dataset. This is because of the even cluster distribution (Verified in Section 5.5), which means approximately $\frac{1}{c}$ of the dataset needs to be visited. Specifically, $O(\frac{nd}{c})$ is the computational cost of Equation 3. $O(\frac{2n}{c})$ is the computational cost of Equation 5, and $O(d)$ is the computational cost of Equation 6.
The space complexity of LIST is $O((l-1)d+dc)+n(d+2)$. $O((l-1)d+dc)$ is the size of the MLP used by our index. $O(nd)$ represents the space required for storing pre-computed $o.emb$, and $O(2n)$ denotes the storage cost for geo-locations.
@@ 5 Experiments
Beijing
Shanghai
Geo-Glue
Number of Pois
122,420
116,859
2,849,754
Number of Queries
168,998
127,183
90,000
Number of Records
233,343
182,634
90,000
Number of Train Queries
136,890
103,019
50,000
Number of Val Queries
15,209
11,446
20,000
Number of Test Queries
16,899
12,718
20,000
Number of Train Records
189,027
148,017
50,000
Number of Val Records
21,034
16,492
20,000
Number of Test Records
23,282
18,125
20,000
We evaluate the effectiveness and efficiency of our proposed solution in answering TkQs by comparing with state-of-the-art methods on three real-world datasets. We aim to answer the following research questions:
  • RQ1: Does our proposed retriever LIST and our proposed relevance model outperform previous state-of-the-art methods in terms of effectiveness?
  • RQ2: How well does the efficiency of our proposed retriever LIST compare with state-of-the-art methods?
  • RQ3: How does LIST scale with the size of dataset?
@@ 5.1 Experimental Setup
Datasets. To evaluate the effectiveness and efficiency of our proposed retriever LIST, we utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue. The Beijing and Shanghai datasets liu2023effectiveness are provided by Meituan, a Chinese retail services platform. Users submit a query through the platform, which consists of a query location and a set of keywords. Subsequently, the Points of Interest (POIs) that users clicked on are recorded in the query log, and is considered as a ground-truth. Note that the click-through data recorded in the search log may be the only feasible way to get a large scale of ground truth data for spatial keyword queries liu2023effectiveness . The explicit feedback such as ratings is very challenging to collect DBLP:conf/www/HeLZNHC17 . Note that using query logs is also the popular way of generating ground truth in the Information Retrieval literature DBLP:conf/www/HeLZNHC17 ; DBLP:conf/cikm/HuangHGDAH13 ; DBLP:conf/www/YaoTCYXD021 ; DBLP:conf/kdd/Joachims02 ; DBLP:journals/csur/WuSZXC23 . Therefore, following the previous work liu2023effectiveness , we utilize the two datasets to evaluate our proposed solution LIST and treat the clicked POIs as ground truth relevant objects to the corresponding queries. The Geo-Glue dataset li2023geoglue ; DBLP:conf/sigir/DingCXHLZX23 is released by DAMO Academic. The POIs in the Geo-Glue dataset are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts. Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
The release of query log datasets from the industry is highly restricted. As a result, to the best of our knowledge, there are no other public datasets that contain ground-truth query results or query logs, except for the three datasets used in our experiments. To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs (shown in Section 5.4)
@@
TkQ
0.5740
0.5283
0.4111
0.3302
0.6746
0.6380
0.5044
0.4009
0.5423
0.5023
0.3847
0.3051
PALM
0.3514
0.3098
0.2077
0.1343
0.4617
0.4023
0.2065
0.1223
N/A
N/A
N/A
N/A
DrW
0.6968
0.6316
0.4814
0.3791
0.7689
0.7159
0.5394
0.4114
N/A
N/A
N/A
N/A
DPR
0.5078
0.4269
0.2892
0.2209
0.5143
0.4161
0.2505
0.1721
0.7626
0.7249
0.6014
0.5079
${}_{D}$
0.7555 0.6845 0.5181 0.4215 0.7986 0.7494 0.5769 0.4568 0.7963 0.7565 0.6211
0.5202
MGeo
N/A
N/A
N/A
N/A
N/A
N/A
N/A
N/A
0.7049
N/A
N/A
0.5270
LIST-R
0.8156 0.7545 0.5913 0.4989 0.8361 0.7924 0.6445 0.5397 0.8393 0.8033 0.6837 0.5887
7.95%
10.22%
14.12%
18.36%
4.69%
5.73%
11.71%
18.14%
5.73%
6.18%
6.26%
11.70%
@@
Baselines of Relevance Model. We compare our proposed relevance model, denoted as LIST-R, with various state-of-the-art spatio-textual relevance models or state-of-the-art deep textual relevance models. We also modify the state-of-the-art textual relevance model DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 into ${}_{D}$ to answer TkQs.
[…]
@@
  • DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 : This is a state-of-the-art textual relevance model for document retrieval. This method does not consider spatial relevance.
  • ${}_{D}$: This method replaces the text relevance computation in DrW liu2023effectiveness with DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 .
@@
  • MGeo DBLP:conf/sigir/DingCXHLZX23 : This is a recent deep relevance based method for answering TkQs. Note that we cannot reproduce the experimental results reported in the paper by running the official code, and we use the evaluation results from the original paper.
[…]
An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by ${}_{D}$. Specifically, on the Beijing and Shanghai datasets, for a fair comparison, we set $k$ to be 10,000, which is the average number of objects retrieved by our index per query. On the geo-glue dataset, $k$ is set to be 55,000 for the same reason above. We evaluate the effect of $k$ in Section 5.3.
@@
  • TkQ+${}_{D}$: This method employs TkQ to retrieve top-$k$ objects and then reranks these objects by ${}_{D}$.
@@
  • IVF+${}_{D}$: This method constructs an IVF index over the embeddings produced by ${}_{D}$, and the objects within the selected index cluster are reranked by ${}_{D}$. Notably, this method requires two parameters: the number of clusters $c$, and the number of clusters to route for each query and object $cr$. It does not involve $k$. We set $c$ and $cr$ to be the same as our method across three datasets for a fair comparison.
[…]
@@
  • LSH+${}_{D}$: This method constructs an LSH index with the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by ${}_{D}$. Following previous work johnson2019billion , we set the length of the hash code $nbits$ to 128.
@@
  • HNSW+${}_{D}$: This method constructs an HNSW index over the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by ${}_{D}$. Following previous work malkovEfficientRobustApproximate2020 , we set the number of links $M$ to 48 and $efConstruction$ to 100.

### v3

@@ 3.2 Data Analysis and Motivations
Motivations of LIST. As discussed in the last paragraph, incorporating deep textual relevance into spatial keyword queries has presented significant efficiency challenges. However, as discussed in Section 2, the first category of deep text relevance model, namely word interaction based deep relevance models, such as DrW liu2023effectiveness , which rely on word interaction modules and word embeddings to compute textual relevance, are not efficient for online computation and retrieval by a specifically designed index. Therefore, we consider embeddings based spatial keyword queries, which capture the semantic meaning of query keywords and object descriptions into two separate embeddings. Then the challenge lies in how to develop an index for embedding based spatial keyword queries as discussed in Introduction.
Spatial Relevance. Previous studies, such as TkQ congEfficientRetrievalTopk2009 and DrW liu2023effectiveness , typically compute spatial relevance by $1-SDist(q.loc,o.loc)$. The implicit assumption behind this linear function is that the user’s geographical preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and we found it does not hold on our real-life datasets. In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$. This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario. This motivates us to design a new spatial relevance module to better fit the real-world scenario. An intuitive solution would be to employ an exponential function. However, our experiments indicate that this method is even less effective compared to the linear function (as to be shown in Section 5.6).
Problem Statement. We aim to develop an ANNS index for deep learning based relevance models designed for embedding based spatial keyword queries, where query keywords and object descriptions are represented as two separate embeddings.
@@ 4 Proposed Retriever (LIST)
Figure 2: The three phases of our retriever LIST: the training, indexing, and query phase. The relevance model is shown in yellow and the index is shown in green.
@@ 4.1 Overview
LIST is featured with a relevance model and a machine learning based ANNS index. Notably, our proposed index is applicable to any other relevance model developed for embedding based spatial keyword queries.
The workflow of LIST is summarized in Figure 2, which has three phases: training, indexing, and query phase. During the training phase, we first train the relevance model, and then train our index. During the indexing phase, each object is partitioned into one cluster. During the query phase, the trained index routes each query to either a single cluster or a subset of clusters that have the highest probabilities. Within these clusters, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
4.2 The Relevance Model
Textual Relevance Learning. Inspired by the dual-encoder model’s success in document retrieval DBLP:conf/emnlp/KarpukhinOMLWEC20 , we employ a dual-encoder module to encode query keywords and object description into two separate embeddings, and calculate textual relevance by the inner product between the two embeddings. Compared with previous relevance models liu2023effectiveness that rely on word embeddings and complex word interaction functions discussed in Section 2, the dual encoder module is both efficient and effective.
The dual-encoder module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, each of which is a BERT model devlin2018bert . The encoder takes the textual content of the query or object as input, captures interactions between words by the transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional embedding. This process is formulated as:
@@ 4.2 The Relevance Model
As discovered in Section 3.2, the user’s geographic preferences for geo-textual objects do not follow a linear pattern. Here, we consider two features of the preference: (1) it increases as the distance decreases, and (2) it exhibits a stepwise decline as the distance increases. The first feature is straightforward because users tend to visit nearby objects. Then a straightforward solution is to design a monotonically continuous function to learn from query-object positive pairs and predict users’ spatial preferences. However, this approach leads to overfitting. For example, if most positive pairs within training dataset are distributed within 1km, the continuous function will learn 1km as a boundary, causing high spatial relevance within 1km but a sharp drop beyond it. Actually, 1.1km is not significantly different from 1km for users. This overfitting results in poor performance due to lack of generalization (as to be shown in Section 5.6). To alleviate this issue, we propose the second feature. To explain this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks. If the nearest Starbucks is very close, s/he would go and buy coffee. If the nearest Starbucks is a little far, s/he may hesitate and her/his intention of visiting the Starbucks would decrease. If the nearest Starbucks is very far away, s/he would give up this idea. Therefore, this pattern aligns with real-world characteristics and alleviates the overfitting issue as the boundaries are manually based on our knowledge. In summary, the new spatial relevance module is designed to be a learnable monotonically step function.
@@
Adaptive Weight Learning. The recent study liu2023effectiveness has shown the importance of weight learning in improving ranking effectiveness. To enhance ranking effectiveness, we propose an adaptive weight learning module, which aims to assign adaptive weights to textual and spatial relevance based on the query keywords. For instance, in the case of detailed query keywords, such as ‘gas station nearby’, giving lower weight to textual relevance better fits the real-world scenario. Since there may be many gas stations nearby, users tend to prefer the closest one. Conversely, in another scenario, such as searching with keywords ‘Lincoln Memorial in Washington, D.C.’, due to the uniqueness of the keyword, giving greater weight to textual relevance can deliver better results.
Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine the weights based on the embedding of query keywords $q.emb$, which is formulated as follows:
@@
Training Strategy. To train our relevance model, we employ the contrastive learning strategy. Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section 5.1). As for the negative (irrelevant) geo-textual objects, previous studies of passage retrieval DBLP:conf/naacl/QuDLLRZDWW21 ; renRocketQAv2JointTraining2021 have shown that hard negative objects can improve retrieval performance. Inspired by them, we introduce this mechanism for spatial keyword query, choosing a subset of hard negative objects for training. Specifically, for each training query, we first filter out the positive objects and then use TkQ congEfficientRetrievalTopk2009 to retrieve a set of top-ranked objects, which are considered the hard negative set for that query. In each training epoch, we randomly pick $b$ hard negative samples from this set. We optimize the loss function as the negative log-likelihood of the positive object:
@@
Complexity Analysis. Now we analyze the time complexity of the proposed relevance model. Assuming the dimensionality of the embedding is $d$, the total number of objects is $n$, and the embeddings of objects are generated in advance. The time complexity of a brute-force search of our relevance model over the entire dataset is $O(n(d+2)+d)$. Here, $O(nd)$ represents the time cost of Equation 3, $O(2n)$ corresponds to the time cost of Equation 5, and $O(d)$ accounts for the time cost of Equation 6.
When compared with other relevance models like DrW liu2023effectiveness , which has a time complexity of $O(nl^{q}(l^{o}d+k^{\prime})+l^{q}(2d^{2}+d))$, our relevance model exhibits superior efficiency in terms of time complexity. Here, $l^{q}$ represents the number of words in the query keywords, $l^{o}$ represents the number of words in the object description, and $k^{\prime}$ denotes the number of top-k’ relevant terms to find in the object description for each query keyword. These hyperparameters, $l^{q}$, $l^{o}$, and $k^{\prime}$, are typically in the dozens. Consequently, DrW is significantly slower than our relevance model.
@@ 4.3 The Proposed Index
Feature Construction. We utilize geo-location and textual embedding to construct a consistent input representation for both objects and queries, which is illustrated in Figure 3. The textual embedding $emb\in\mathbb{R}^{d}$ is a $d$-dimensional embedding converted from textual content $o.doc$ or $q.doc$ by the trained dual encoder module, which is typically hunderds, e.g., 768. It is L2 normalized before being input into the neural network. The geo-location $loc=\langle lat,lon\rangle$ are transformed into the following features $\langle\hat{lat},\hat{lon}\rangle$ as below:
@@
Here $x$ is the input feature vector for the cluster classifier, i.e., the MLP. Since the two feature vectors are normalized independently, they will have a similar impact on the neural network.
@@
Insertion and Deletion Policy. When a new object comes, we convert it into an embedding using the trained relevance model and then assign it to specific clusters using the trained index. When an object is deleted, we simply remove it from the corresponding cluster. The time cost of the insertion operation is equivalent to the inference time of the neural networks, while the time cost of the deletion operation involves scanning the cluster lists and then deleting the object’s id. The time costs of both operations are negligible.
Note that most existing ANNS indexes are static luoSurveyDeepHashing2023 ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 . When inserted data significantly differs in distribution from the existing data or when insertions occur frequently, periodically rebuilding the index is necessary to maintain high accuracy. This is because most ANNS indexes are based on the principle of organizing similar objects together, thereby routing queries to similar blocks to reduce search space. The functions determining which cluster each object belongs to are typically static, such as the distance between objects and the centroids of each cluster in an IVF index. When the distribution of embeddings changes significantly, these static functions can no longer accurately capture the new distribution, leading to a drop in performance. In our approach, we address this issue by only retraining the index, but not the relevance model.
@@ 18       $Res_{q}\leftarrow\text{argTop-k}_{o\in C_{i}}ST(q,o)$;
Procedures of LIST. The detailed procedures of LIST are summarized at Algorithm 1. In the training phase, we train the proposed relevance model and index (lines 2-4). After that, in the indexing phase, we assign all objects to their corresponding clusters as inverted files (lines 6-10). In the query phase, given a new query $q$, we extract its features and then route it to a cluster $C_{i}$ by the index (lines 13-16). Consequently, we calculate the relevance score between $q$ and each object $o$ in that cluster by the relevance model, and the top-$k$ objects with the highest scores are returned to answer query $q$ (lines 17-18). In practice, the number of objects assigned to each cluster is relatively small, and thus we evaluate all objects within the routed clusters as it is in existing ANNS indexes.
Complexity Analysis. Now we analyze the time and space complexity of LIST. Assuming the dimensionality of the embedding is $d$, the total number of objects is $n$, the number of clusters is $c$, and the number of layers in the MLP of the index is $l$. The embeddings $o.emb$ of objects are generated in advance.
For a given query, the time complexity of LIST is $O((l-2)d^{2}+dc+\frac{n}{c}(d+2)+d)$. $O((l-2)d^{2}+dc)$ is the time complexity for our index (lines 15), which is the inference computation cost of Equation 11. $O(\frac{n}{c}(d+2)+d)$ is the time cost of our relevance model (line 17). $\frac{n}{c}$ denotes the number of objects to be checked by the relevance model, which is approximately $\frac{1}{c}$ of the entire dataset. This is because of the even cluster distribution (Verified in Section 5.6), which means approximately $\frac{1}{c}$ of the dataset needs to be visited. The space complexity of LIST is $O((l-1)d+dc)+n(d+2)$. $O((l-1)d+dc)$ is the size of the MLP used by our index. $O(nd)$ represents the space required for storing pre-computed $o.emb$, and $O(2n)$ denotes the storage cost for geo-locations.
@@ 5 Experiments
In this section, we evaluate the effectiveness and efficiency of our proposed solution for answering TkQs by comparing it with state-of-the-art methods on three real-world datasets. We aim to answer the following research questions:
  • RQ1: Does our relevance model outperform existing relevance models in terms of effectiveness?
  • RQ2: Does our proposed ANNS index achieve a better effectiveness-efficiency trade-off compared to existing indexes?
  • RQ3: Can our proposed ANNS index be applied to other relevance models designed for embedding based spatial keyword queries?
  • RQ4: How does LIST scale with the size of dataset?
@@
Beijing Shanghai Geo-Glue
Number of Pois 122,420 116,859 2,849,754
Number of Queries 168,998 127,183 90,000
Number of Records 233,343 182,634 90,000
Number of Train Queries 136,890 103,019 50,000
Number of Val Queries 15,209 11,446 20,000
Number of Test Queries 16,899 12,718 20,000
Number of Train Records 189,027 148,017 50,000
Number of Val Records 21,034 16,492 20,000
Number of Test Records 23,282 18,125 20,000
@@ 5.1 Experimental Setup
Datasets. To evaluate the effectiveness and efficiency of our proposed retriever LIST, we utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue. Among them, the Beijing and Shanghai datasets liu2023effectiveness are provided a Chinese retail services platform. Users submit a query through the platform, which consists of a query location and a set of keywords. Subsequently, the Points of Interest (POIs) that users clicked on are recorded in the query log, and is considered as a ground-truth. Note that the click-through data recorded in the search log may be the only feasible way to get a large scale of ground truth data for spatial keyword queries liu2023effectiveness . The explicit feedback such as ratings is very challenging to collect DBLP:conf/www/HeLZNHC17 . Note that using query logs is also the popular way of generating ground truth in the Information Retrieval literature DBLP:conf/www/HeLZNHC17 ; DBLP:conf/cikm/HuangHGDAH13 ; DBLP:conf/www/YaoTCYXD021 ; DBLP:conf/kdd/Joachims02 . Therefore, following the previous work liu2023effectiveness , we utilize the two datasets to evaluate our proposed solution LIST and treat the clicked POIs as ground truth relevant objects to the corresponding queries. In the Geo-Glue dataset li2023geoglue ; DBLP:conf/sigir/DingCXHLZX23 . the POIs are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts. Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
The release of query log datasets from the industry is highly restricted. As a result, to the best of our knowledge, there are no other public datasets that contain ground-truth query results or query logs, except for the three datasets used in our experiments. To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs from OpenStreetMap (as to be shown in Section 5.5)
Table 2: Hyperparameter settings on three datasets.
Hyperparameters Beijing Shanghai Geo-Glue
spatial footstep $t$ 1000 1000 1000
pseudo-label’s $\text{neg}_{\text{start}}$ 50,000 60,000 180,000
pseudo-label’s $\text{neg}_{\text{end}}$ 55,000 65,000 181,000
number of cluster $c$ 20 20 300
@@
TkQ 0.5740 0.5283 0.4111 0.3302 0.6746 0.6380 0.5044 0.4009 0.5423 0.5023 0.3847 0.3051
PALM 0.3514 0.3098 0.2077 0.1343 0.4617 0.4023 0.2065 0.1223 N/A N/A N/A N/A
DrW 0.6968 0.6316 0.4814 0.3791 0.7689 0.7159 0.5394 0.4114 N/A N/A N/A N/A
MGeo N/A N/A N/A N/A N/A N/A N/A N/A 0.7049 N/A N/A 0.5270
LIST-R 0.8156 0.7545 0.5913 0.4989 0.8361 0.7924 0.6445 0.5397 0.8393 0.8033 0.6837 0.5887
(Gain) 17.04% 19.45% 22.82% 31.60% 10.63% 19.48% 19.06% 18.14% 19.06% 59.92% 77.72% 11.70%
@@
Relevance Model Baselines. We compare our relevance model, denoted as LIST-R, with existing state-of-the-art spatio-textual relevance models.
  • TkQ congEfficientRetrievalTopk2009 : It uses traditional relevance model BM25 DBLP:journals/ftir/RobertsonZ09 to evaluate text relevance and treats $1-SDist(q.loc,o.loc)$ as spatial relevance. The weight parameter $\alpha$ is manually tuned from 0 to 1 with a footstep of 0.1, and the best effectiveness is achieved when $\alpha=0.4$ for all three datasets.
@@
  • MGeo DBLP:conf/sigir/DingCXHLZX23 : This is a recent deep learning based method for answering TkQs. Note that we cannot reproduce the experimental results reported in the paper by running the official code, and thus we use the evaluation results from the original paper.
Index Baselines. We compare our proposed index with existing state-of-the-art indexes. Specifically, we select the IR-Tree congEfficientRetrievalTopk2009 to accelerate the TkQ search, serving as an index baseline. As discussed in Section 2, existing ANNS indexes fall into two categories and our method belongs to the first category. We select the state-of-the-art ANNS indexes in the first category, i.e., IVF jegouProductQuantizationNearest2011 , LSH johnson2019billion , and HNSW malkovEfficientRobustApproximate2020 , as index baselines. For the second category of methods, such as product quantization based methods, their standalone efficiency is no better than a brute-force search [^2] https://github.com/facebookresearch/faiss/issues/148. Therefore, following previous work jegouProductQuantizationNearest2011 , we choose IVFPQ as an index baseline, which integrates PQ with the IVF index from the first category and is considered as a state-of-the-art solution. As discussed in Section 3.2, existing ANNS indexes are limited to considering only textual embeddings, neglecting the spatial factor. To mitigate this limitation, we extend the IVF index to include both embedding and spatial factors, which is denoted as IVFS index, as an index baseline. Similar to LIST, which uses a retrieve and rerank pipeline, we utilize the aforementioned indexes to retrieve a subset of objects and then rerank the retrieved objects with our relevance model LIST-R for a fair comparison. The integration of our proposed index and LIST-R is LIST, and the other methods are detailed below.
@@
  • TkQ+LIST-R: This method employs TkQ to retrieve top-$k$ objects and reranks these objects by LIST-R.
@@
  • IVF+LIST-R: This method constructs an IVF index over the embeddings produced by LIST-R, and the objects within the selected index cluster are reranked by LIST-R. Notably, this method requires two parameters: the number of clusters $c$, and the number of clusters to route for each query and object $cr$. It does not involve $k$. We set $c$ and $cr$ to be the same as our method across three datasets for a fair comparison.
  • IVFS+LIST-R: This method follows the IVF+LIST-R pipeline but diverges in the clustering approach. Instead of only applying k-means to text embeddings, this approach utilizes k-means on the weighted sum of both geo-location and text embedding factors. The weight $\alpha$ is manually tuned from 0 to 1 with a footstep of 0.1, and the best effectiveness is achieved when $\alpha=0.9$ for all these datasets.
@@
  • LSH+LIST-R: This method constructs an LSH index with the embeddings produced by LIST-R, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by LIST-R. Following previous work johnson2019billion , we set the length of the hash code $nbits$ to 128.
@@
  • HNSW+LIST-R: This method constructs an HNSW index over the embeddings produced by LIST-R, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by LIST-R. Following previous work malkovEfficientRobustApproximate2020 , we set the number of links $M$ to 48 and $efConstruction$ to 100.